OpenStudy (anonymous):

Suppose a laboratory has a 57 g sample of polonium-210. The half-life of polonium-210 is about 138 days. a. How many half-lives of polonium-210 occur in 276 days? b. How much polonium is in the sample 276 days later? 2; 3,933 g 2; 14.25 g 2; 28.5 g 3; 7.13 g

OpenStudy (anonymous):

one half life = 138 days, 2 half lives = 276 days, in the first half life half od the polonium decays, in the second half life half od what remained decays: 1/2 * 1/2 = 1/4 = 0.25